OpenStudy (anonymous):

Max has a monthly salary of $1,100 and also earns 7.5% commission on his sales. If Max had $43,000 in sales last month, what were his total earnings for the month? a. $3,225 b. $4,050 c. $4,325 d. $5,150

OpenStudy (anonymous):

total earning = monthly salary + commission = 1100 + 7.5% of 43000 = 1100 + 7.5/100 * 43000 = 1100 + 7.5 * 430 = 1100 +3225 = 4325 dollars Hence option C is the correct answer. @pflanagan23

OpenStudy (anonymous):

yeah thanks